OpenStudy (anonymous):

find answers between cos2x=cosx Stars given!!

OpenStudy (jdoe0001):

\(\bf cos(2x)= cos(x)\\ \textit{so we know that } cos(2x) = 2cos^2(x)-1\\ 2cos^2(x)-1= cos(x) \implies 2cos^2(x)-1-cos(x) = 0\\ 2cos^2(x)-cos(x)-1 = 0\)

OpenStudy (jdoe0001):

so is a quadratic equation, then just solve for cos(x)

OpenStudy (jdoe0001):

cos(x) = a \(\bf 2cos^2(x)-cos(x)-1 = 0 \implies 2a^2-a-1 = 0\)

OpenStudy (jdoe0001):

\(\bf \textit{so we know that } cos(2x) = 2cos^2(x)-1\) so replace your cos(2x) for that

OpenStudy (anonymous):

?

OpenStudy (jdoe0001):

is a trig identity http://www.freemathhelp.com/images/doubleangles.png